# references.bib

### v1

@@ E.2 Step 2: Convergence of component $X^{\ell,1}_{t}$
$h^{\ell}(S_{t},Z_{t},A_{t},Z_{t+1})=\bigl{[}r(S_{t},A_{t})-r^{\ell}_{\mu}(Z_{t},A_{t})\bigr{]}+\Bigl{[}\gamma V^{\llbracket\ell+1\rrbracket}(Z_{t+1})-\gamma\sum_{z^{\prime}\in\mathsf{Z}}P^{\ell}_{\mu}(z^{\prime}|Z_{t},A_{t})V^{\llbracket\ell+1\rrbracket}_{\mu}(z^{\prime})\Bigr{]}.$ (62)
Then, Prop. 4 implies that the sequence $\{X^{\ell,1}_{t}\}_{t\geq 1}$ converges almost surely to the limit
$\sum_{(s,z_{+})\in\mathsf{S}\times\mathsf{Z}}\zeta^{\ell}_{\mu}(s,z,a,z_{+})h^{\ell}(s,z,a,z_{+})$ (63)
where $\zeta^{\ell}_{\mu}(s,z,a,z_{+})=\zeta^{\ell}_{\mu}(s,z,a)P^{\ell}_{\mu}(z_{+}|z,a)$. Furthermore, the Strong Law of Large numbers for periodic Markov chains (Prop. 3, Eq. (37)) implies that the term in (63) converges almost surely to zero.
Hence, for all $(\ell,z,a)$, the process $\{X^{\ell,1}_{t}(z,a)\}_{t\geq 1}$ converges to zero almost surely.
E.3 Step 3: Convergence of component $X^{\ell,2}_{t}$
The remaining analysis is similar to corresponding step in the standard convergence proof of Q-learning and its variations [Jaakkola1994, Kara2022, Kara2024].
@@ E.3 Step 3: Convergence of component $X^{\ell,2}_{t}$
Arbitrarily fix a $\epsilon>0$. In the previous step, we have shown that $\lVert X^{\ell,1}_{t}\rVert\to 0$ a.s. Therefore, there exists a set $\Omega^{1}$ of measure one and a constant $T(\omega,\epsilon)$ such that for $\omega\in\Omega^{1}$, all $t>T(\omega,\epsilon)$, and $(z,a)\in\mathsf{Z}\times\mathsf{A}$, we have
$X^{\ell,1}_{t}(z,a)<\epsilon.$ (64)
@@
Suppose for some $t>T(\omega,\epsilon)$, $\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, i.e., $\epsilon<\lVert X^{\ell,2}\rVert/C$. Then, for any $(z,a)\in\mathsf{Z}\times\mathsf{A}$,
$U^{\ell,2}_{t}(z,a)$ $\leq\gamma\lVert V^{\ell}_{t}-V^{\ell}_{\mu}\rVert\leq\gamma\lVert Q^{\ell}_{t}-Q^{\ell}_{\mu}\rVert=\gamma\lVert\Delta_{t}\rVert$ (66a)
$\leq\gamma\lVert X^{\ell,1}_{t}\rVert+\gamma\lVert X^{\ell,2}_{t}\rVert\leq\gamma\epsilon+\gamma\lVert X^{\ell,2}_{t}\rVert$ (66b)
$\leq\gamma\left(1+\frac{1}{C}\right)\lVert X^{\ell,2}_{t}\rVert=\kappa\lVert X^{\ell,2}_{t}\rVert<\lVert X^{\ell,2}_{t}\rVert.$ (66c)
Thus, for any $t>T(\omega,\epsilon)$ and $\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, we have
$X^{\ell,2}_{t+1}(z,a)=(1-\alpha^{\ell}_{t}(z,a))X^{\ell,2}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)U^{\ell,2}_{t}(z,a)<\lVert X^{\ell,2}_{t}\rVert.$ (67)
Hence, when $\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, it decreases monotonically. Hence, there are two possibilities: either (i) $\lVert X^{\ell,2}_{t}\rVert$always remains above $C\epsilon$; or (ii) it goes below $C\epsilon$ at some stage. We consider these two possibilities separately.
E.3.1 Possibility (i): $\lVert X^{\ell,2}_{t}\rVert$ always remains above $C\epsilon$
We will show that $\lVert X^{\ell,2}_{t}\rVert$ cannot remain above $C\epsilon$ for ever. We first start with a basic result for random iterations. This is a self-contained result, so we reuse some of the variables used in the rest of the paper.
@@ E.3.1 Possibility (i): $\lVert X^{\ell,2}_{t}\rVert$ always remains above $C\epsilon$
We will now prove that $\lVert X^{\ell,2}_{t}\rVert$ cannot remain above $C\epsilon$ forever. The proof is by contraction. Suppose $\lVert X^{\ell,2}_{t}\rVert$ remains above $C\epsilon$ forever. As argued earlier, this implies that $\lVert X^{\ell,2}_{t}\rVert$, $t\geq T(\omega,\epsilon)$ is a strictly decreasing sequence, so it must be bounded from above. Let $B^{(0)}$ be such that $\lVert X^{\ell,2}_{t}\rVert\leq B^{(0)}$ for all $t\geq T(\omega,\epsilon)$. Eq. (66c) implies that $\lVert U^{\ell,2}_{t}\rVert\leq\kappa B^{(0)}$. Thus,
$X^{\ell,2}_{t+1}(z,a)\leq(1-\alpha^{\ell}_{t}(z,a))\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\lVert U^{\ell,2}_{t}\rVert\leq(1-\alpha^{\ell}_{t}(z,a))\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\kappa\lVert X^{\ell,2}_{t}\rVert$ (72)
which implies that $\lVert X^{\ell,2}_{t}\rVert\leq\lVert M^{\ell,(0)}_{t}\rVert$, where $\{M^{\ell,(0)}_{t}\}_{t\geq T(\omega,\epsilon)}$ is a sequence given by
@@
Lem. 2 implies that $M^{\ell,(0)}_{t}(z,a)\to\kappa B^{(0)}$ and hence $\lVert M^{\ell,(0)}_{t}\rVert\to\kappa B^{(0)}$. Now pick an arbitrary $\bar{\epsilon}\in(0,(1-\kappa)C\epsilon)$. Thus, there exists a time $T^{(1)}=T^{(1)}(\omega,\epsilon,\bar{\epsilon})$ such that for all $t>T^{(1)}$, $\lVert M^{\ell,(0)}_{t}\rVert\leq B^{(1)}\coloneqq\kappa B^{(0)}+\epsilon$. Since $\lVert X^{\ell,2}_{t}\rVert$ is bounded by $\lVert M^{\ell,(0)}_{t}\rVert$, this implies that for all $t>T^{(1)}$, $\lVert X^{\ell,2}_{t}\rVert\leq B^{(1)}$ and, by (66c), $\lVert U^{\ell,2}_{t}\rVert\leq\kappa B^{(1)}$. By repeating the above argument, there exists a time $T^{(2)}$ such that for all $t\geq T^{(2)}$,
$\lVert X^{\ell,2}_{t}\rVert\leq B^{(2)}\coloneqq\kappa B^{(1)}+\bar{\epsilon}=\kappa^{2}B^{(0)}+\kappa\bar{\epsilon}+\bar{\epsilon},$ (74)
and so on. By (65), $\kappa<1$ and $\bar{\epsilon}$ is chosen to be less than $C\epsilon$. So eventually, $B^{(m)}\coloneqq\kappa^{m}B^{(0)}+\kappa^{m-1}\bar{\epsilon}+\cdots+\bar{\epsilon}$ must get below $C\epsilon$ for some $m$, contradicting the assumption that $\lVert X^{\ell,2}_{t}\rVert$ remains above $C\epsilon$ forever.
E.3.2 Possibility (ii): $\lVert X^{\ell,2}_{t}\rVert$ goes below $C\epsilon$ at some stage
Suppose that there is some $t>T(\omega,\epsilon)$ such that $\lVert X^{\ell,2}_{t}\rVert<C\epsilon$. Then (66b) implies that
$\lVert U^{\ell,2}_{t}\rVert\leq\gamma\lVert X^{\ell,1}_{t}\rVert+\gamma\lVert X^{\ell,2}_{t}\rVert\leq\gamma\epsilon+\gamma C\epsilon<C\epsilon$ (75)
@@ E.3.2 Possibility (ii): $\lVert X^{\ell,2}_{t}\rVert$ goes below $C\epsilon$ at some stage
$X^{\ell,2}_{t+1}(z,a)\leq(1-\alpha^{\ell}_{t}(z,a))\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\lVert U^{\ell,2}_{t}\rVert<C\epsilon$ (76)
where the last inequality uses the fact that both $\lVert U^{\ell,2}_{t}\rVert$ and $\lVert X^{\ell,2}_{t}\rVert$ are both below $C\epsilon$. Thus, we have that
$\lVert X^{\ell,2}_{t}\rVert<C\epsilon.$ (77)
Hence, once $\lVert X^{\ell,2}_{t}\rVert$ goes below $C\epsilon$, it stays there.
E.3.3 Implication
We have show that for sufficiently large $t>T(\omega,\epsilon)$, $\lVert X^{\ell,2}_{t}\rVert<C\epsilon$. Since $\epsilon$ is arbitrary, this means that for all $\omega\in\Omega^{1}$, $\lVert X^{\ell,2}_{t}\rVert\to 0$. Thus,
$\lim_{t\to\infty}\lVert X^{\ell,2}_{t}\rVert=0,\quad a.s.$ (78)
E.4 Putting everything together
Recall that we defined $\Delta^{\ell}_{t}=Q^{\ell}_{t}-Q_{\mu}$ and split $\Delta^{\ell}_{t}=X^{\ell,1}_{t}+X^{\ell,2}_{t}$. Step 2 shows that $\lVert X^{\ell,1}_{t}\rVert\to 0$, a.s. and Eq. (78) shows that $\lVert X^{\ell,2}_{t}\rVert\to 0$, a.s. Thus, by triangle inequality,
$\lim_{t\to\infty}\lVert\Delta^{\ell}_{t}\rVert\leq\lim_{t\to\infty}\lVert X^{\ell,1}_{t}\rVert+\lim_{t\to\infty}\lVert X^{\ell,2}_{t}\rVert=0,$ (79)

### v3

@@ E.3 Step 3: Convergence of component $X^{\ell,1}_{t}$
$h_{P}(W_{t};\ell,z_{\circ},a_{\circ})=\Bigl{[}\gamma V^{\llbracket\ell+1\rrbracket}_{\mu}(Z_{t+1})-\gamma\sum_{\bar{z}\in\mathsf{Z}}P^{\ell}_{\mu}(\bar{z}|z_{\circ},a_{\circ})V^{\llbracket\ell+1\rrbracket}_{\mu}(\bar{z})\Bigr{]}\mathds{1}_{\{Z_{t}=z_{\circ},A_{t}=a_{\circ}\}}$ (64)
Then the process $\{X^{\ell,1}_{t}(z,a)\}_{t\geq 1}$ is given by the stochastic iteration
$X^{\ell,1}_{t+1}(z_{\circ},a_{\circ})=(1-\alpha^{\ell}_{t}(z_{\circ},a_{\circ}))X^{\ell,1}_{t}(z_{\circ},a_{\circ})+\alpha^{\ell}_{t}(z_{\circ},a_{\circ})h_{P}(W_{t};\ell,z_{\circ},a_{\circ}).$ (65)
which is of the form (56). As argued earlier, the process $\{W_{t}\}_{t\geq 1}$ is a periodic Markov chain. Due to Assm. 1, the learning rate $\alpha^{\ell}_{t}(z_{\circ},a_{\circ})$ is measurable with respect to the sigma-algebra generated by $(Z_{1:t},A_{1:t})$ and is therefore also measurable with respect to the sigma-algebra generated by $W_{1:t}$. Combining this with Prop. 4 implies that the learning rates $\{\alpha^{\ell}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ satisfy the conditions of Prop. 4. Therefore, Prop. 4 implies that $\{X^{\ell,1}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ converges a.s. to the following limit
$\hskip-10.00002pt\lim_{t\to\infty}X^{\ell,1}_{t}(z_{\circ},a_{\circ})$
$=\sum_{\begin{subarray}{c}s,z,a\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}\\ s^{\prime},z^{\prime},a^{\prime}\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}\end{subarray}}\bar{\zeta}^{\ell}_{\mu}(s,z,a,s^{\prime},z^{\prime},a^{\prime})h_{P}(s,z,a,s^{\prime},z^{\prime},a^{\prime};\ell,z_{\circ},a_{\circ})$
$=\sum_{\begin{subarray}{c}s,z,a\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}\\ s^{\prime},z^{\prime},a^{\prime}\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}\end{subarray}}\bar{\zeta}^{\ell}_{\mu}(s,z,a,s^{\prime},z^{\prime},a^{\prime})\Bigl{[}\gamma V^{\llbracket\ell+1\rrbracket}_{\mu}(z^{\prime})-\gamma\sum_{\bar{z}\in\mathsf{Z}}P^{\ell}_{\mu}(\bar{z}|z_{\circ},a_{\circ})V^{\llbracket\ell+1\rrbracket}_{\mu}(\bar{z})\Bigr{]}\mathds{1}_{\{z=z_{\circ},a=a_{\circ}\}}$
$=\gamma\biggl{[}\sum_{z^{\prime}\in\mathsf{Z}}\bar{\zeta}^{\ell}_{\mu}(\mathsf{S},z_{\circ},a_{\circ},\mathsf{S},z^{\prime},\mathsf{A})V^{\llbracket\ell+1\rrbracket}_{\mu}(z^{\prime})\biggr{]}-\biggl{[}\gamma\bar{\zeta}^{\ell}_{\mu}(\mathsf{S},z_{\circ},a_{\circ},\mathsf{S},\mathcal{Z},\mathcal{A})\sum_{\bar{z}\in\mathsf{Z}}P^{\ell}_{\mu}(\bar{z}|z_{\circ},a_{\circ})V^{\llbracket\ell+1\rrbracket}_{\mu}(\bar{z})\biggr{]}$
$=0$
where the last step follows from the fact that $\bar{\zeta}^{\ell}_{\mu}(\mathsf{S},z_{\circ},a_{\circ},\mathsf{S},\mathcal{Z},\mathcal{A})=\zeta_{\mu}^{\ell}(z_{\circ},a_{\circ})$ and $\bar{\zeta}^{\ell}_{\mu}(\mathsf{S},z_{\circ},a_{\circ},\mathsf{S},z^{\prime},\mathsf{A})=\zeta_{\mu}^{\ell}(z_{\circ},a_{\circ})P^{\ell}_{\mu}(z^{\prime}|z_{\circ},a_{\circ})$.
E.4 Step 4: Convergence of component $X^{\ell,2}_{t}$
[…]
@@ E.4 Step 4: Convergence of component $X^{\ell,2}_{t}$
In the previous step, we have shown that $\lVert X^{\ell,i}_{t}\rVert\to 0$ a.s., for $i\in\{0,1\}$. Thus, we have that $\lVert X^{\ell,0}_{t}+X^{\ell,1}_{t}\rVert\to 0$ a.s. Arbitrarily fix an $\epsilon>0$. Therefore, there exists a set $\Omega^{1}$ of measure one and a constant $T(\omega,\epsilon)$ such that for $\omega\in\Omega^{1}$, all $t>T(\omega,\epsilon)$, and $(\ell,z,a)\in\mathsf{L}\times\mathsf{Z}\times\mathsf{A}$, we have
$X^{\ell,0}_{t}(z,a)+X^{\ell,1}_{t}(z,a)<\epsilon.$ (66)
@@
Suppose for some $t>T(\omega,\epsilon)$, $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert>C\epsilon$. Then, for $(z,a)\in\mathsf{Z}\times\mathsf{A}$,
$U^{\ell,2}_{t}(z,a)$ $=\gamma V_{t}^{\llbracket\ell+1\rrbracket}(Z_{t+1})-\gamma V^{\llbracket\ell+1\rrbracket}_{\mu}(Z_{t+1})$ (68a)
$=\gamma\max_{a\in\mathsf{A}}Q_{t}^{\llbracket\ell+1\rrbracket}(Z_{t+1},a)-\max_{a^{\prime}\in\mathsf{A}}\gamma Q^{\llbracket\ell+1\rrbracket}_{\mu}(Z_{t+1},a^{\prime})$ (68b)
$\leq\gamma\max_{a\in\mathsf{A}}\left\{Q_{t}^{\llbracket\ell+1\rrbracket}(Z_{t+1},a)-\gamma Q^{\llbracket\ell+1\rrbracket}_{\mu}(Z_{t+1},a)\right\}$ (68c)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\gamma\lVert Q^{\llbracket\ell+1\rrbracket}_{t}-Q^{\llbracket\ell+1\rrbracket}_{\mu}\rVert=\gamma\lVert\Delta^{\llbracket\ell+1\rrbracket}_{t}\rVert$ (68d)
$\leq\gamma\lVert X^{\llbracket\ell+1\rrbracket,0}_{t}+X^{\llbracket\ell+1\rrbracket,1}_{t}\rVert+\gamma\lVert X^{\llbracket\ell+1\rrbracket,2}_{t}\rVert\stackrel{{\scriptstyle(b)}}{{\leq}}\gamma\epsilon+\gamma\lVert X^{\llbracket\ell+1\rrbracket,2}_{t}\rVert$ (68e)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\gamma\left(1+\frac{1}{C}\right)\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\stackrel{{\scriptstyle(d)}}{{=}}\kappa\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\stackrel{{\scriptstyle(d)}}{{<}}\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert.$ (68f)
where $(a)$ follows from the fact that an upper bound is obtained by maximizing over all realizations of $Z_{t+1}$, $(b)$ follows from (66), $(c)$ follows from the fact that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, $(d)$ follows from (67). Thus, for any $t>T(\omega,\epsilon)$ and $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, we have
$X^{\ell,2}_{t+1}(z,a)$ $=(1-\alpha^{\ell}_{t}(z,a))X^{\ell,2}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)U^{\ell,2}_{t}(z,a)<\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$
$\implies\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t+1}\rVert$ $<\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert.$
Hence, when $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert>C\epsilon$, it decreases monotonically with time. Hence, there are two possibilities: either (i) $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$always remains above $C\epsilon$; or (ii) it goes below $C\epsilon$ at some stage. We consider these two possibilities separately.
E.4.1 Possibility (i): $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ always remains above $C\epsilon$
We will show that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ cannot remain above $C\epsilon$ forever. We first start with a basic result for random iterations. This is a self-contained result, so we reuse some of the variables used in the rest of the paper.
@@
We will now prove that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ cannot remain above $C\epsilon$ forever. The proof is by contradiction. Suppose $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ remains above $C\epsilon$ forever. As argued earlier, this implies that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$, $t\geq T(\omega,\epsilon)$, is a strictly decreasing sequence, so it must be bounded from above. Let $B^{(0)}$ be such that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\leq B^{(0)}$ for all $t\geq T(\omega,\epsilon)$. Eq. (68f) implies that $\lVert U^{\ell,2}_{t}\rVert<\kappa B^{(0)}$. Then, we have that
$\max_{\ell\in\mathsf{L}}X^{\ell,2}_{t+1}(z,a)$ $\leq(1-\alpha^{\ell}_{t}(z,a))\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\max_{\ell\in\mathsf{L}}\lVert U^{\ell,2}_{t}\rVert$
$\leq(1-\alpha^{\ell}_{t}(z,a))\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\kappa\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$
which implies that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\leq\lVert M^{\ell,(0)}_{t}\rVert$, where $\{M^{\ell,(0)}_{t}\}_{t\geq T(\omega,\epsilon)}$ is a sequence given by
@@
Lem. 2 implies that $M^{\ell,(0)}_{t}(z,a)\to\kappa B^{(0)}$ and hence $\lVert M^{\ell,(0)}_{t}\rVert\to\kappa B^{(0)}$. Now pick an arbitrary $\bar{\epsilon}\in(0,(1-\kappa)C\epsilon)$. Thus, there exists a time $T^{(1)}=T^{(1)}(\omega,\epsilon,\bar{\epsilon})$ such that for all $t>T^{(1)}$, $\lVert M^{\ell,(0)}_{t}\rVert\leq B^{(1)}\coloneqq\kappa B^{(0)}+\bar{\epsilon}$. Since $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ is bounded by $\lVert M^{\ell,(0)}_{t}\rVert$, this implies that for all $t>T^{(1)}$, $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\leq B^{(1)}$ and, by (68f), $\lVert U^{\ell,2}_{t}\rVert\leq\kappa B^{(1)}$. By repeating the above argument, there exists a time $T^{(2)}$ such that for all $t\geq T^{(2)}$,
$\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\leq B^{(2)}\coloneqq\kappa B^{(1)}+\bar{\epsilon}=\kappa^{2}B^{(0)}+\kappa\bar{\epsilon}+\bar{\epsilon},$ (74)
and so on. By (67), $\kappa<1$ and $\bar{\epsilon}$ is chosen to be less than $C\epsilon$. So eventually, $B^{(m)}\coloneqq\kappa^{m}B^{(0)}+\kappa^{m-1}\bar{\epsilon}+\cdots+\bar{\epsilon}$ must get below $C\epsilon$ for some $m$, contradicting the assumption that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ remains above $C\epsilon$ forever.
E.4.2 Possibility (ii): $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert$ goes below $C\epsilon$ at some stage
Suppose that there is some $t>T(\omega,\epsilon)$ such that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert<C\epsilon$. Then (68e) implies that
$\lVert U^{\ell,2}_{t}\rVert\leq\gamma\lVert X^{\llbracket\ell+1\rrbracket,0}_{t}+X^{\llbracket\ell+1\rrbracket,1}_{t}\rVert+\gamma\lVert X^{\llbracket\ell+1\rrbracket,2}_{t}\rVert\leq\gamma\epsilon+\gamma C\epsilon<C\epsilon$ (75)
@@
$\max_{\ell\in\mathsf{L}}X^{\ell,2}_{t+1}(z,a)\leq(1-\alpha^{\ell}_{t}(z,a))\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert+\alpha^{\ell}_{t}(z,a)\max_{\ell\in\mathsf{L}}\lVert U^{\ell,2}_{t}\rVert<C\epsilon$ (76)
where the last inequality uses the fact that both $\lVert U^{\ell,2}_{t}\rVert$ and $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t+1}\rVert$ are both below $C\epsilon$. Thus, we have that
$\max_{\ell\in\mathsf{L}}X^{\ell,2}_{t+1}(z,a)<C\epsilon.$ (77)
Hence, once $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t+1}\rVert$ goes below $C\epsilon$, it stays there.
E.4.3 Implication
We have show that for sufficiently large $t>T(\omega,\epsilon)$, $\max_{\ell\in\mathsf{L}}X^{\ell,2}_{t}(z,a)<C\epsilon$. Since $\epsilon$ is arbitrary, this means that for all realizations $\omega\in\Omega^{1}$, $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\to 0$. Thus,
$\lim_{t\to\infty}\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert=0,\quad a.s.$ (78)
E.5 Putting everything together
Recall that we defined $\Delta^{\ell}_{t}=Q^{\ell}_{t}-Q_{\mu}$ and in Step $1$, we split $\Delta^{\ell}_{t}=X^{\ell,0}_{t}+X^{\ell,1}_{t}+X^{\ell,2}_{t}$. Steps $2$ and $3$ together show that $\lVert X^{\ell,0}_{t}+X^{\ell,1}_{t}\rVert\to 0$, a.s. and Step $3$ (78) shows us that $\max_{\ell\in\mathsf{L}}\lVert X^{\ell,2}_{t}\rVert\to 0$, a.s. Thus, by the triangle inequality,
$\lim_{t\to\infty}\lVert\Delta^{\ell}_{t}\rVert\leq\lim_{t\to\infty}\lVert X^{\ell,0}_{t}+X^{\ell,1}_{t}\rVert+\lim_{t\to\infty}\lVert X^{\ell,2}_{t}\rVert=0,$ (79)